OpenStudy (anonymous):

what is the simplest form of (4x^2-25)/(2x-5)

jimthompson5910 (jim_thompson5910):

(4x^2-25)/(2x-5) ((2x-5)(2x+5))/(2x-5) 2x+5

OpenStudy (anonymous):

im still confused . can you explain how you got this?like in words?

jimthompson5910 (jim_thompson5910):

I factored 4x^2-25 to get (2x-5)(2x+5) using the difference of squares rule

jimthompson5910 (jim_thompson5910):

then i divided and canceled out the "2x-5" terms

jimthompson5910 (jim_thompson5910):

to be left with 2x+5 as the answer

OpenStudy (anonymous):

ok but the answers show more. can i show you my choices?

jimthompson5910 (jim_thompson5910):

sure thing

OpenStudy (anonymous):

2x – 5, with the restriction x ≠ – 5/2 2x + 5, with the restriction x ≠ 5/2 2x – 5, with the restriction x ≠ 5/2 2x + 5, with the restriction x ≠ – 5/2

jimthompson5910 (jim_thompson5910):

well you can't divide by zero, so 2x-5 cannot be zero if it were, then 2x-5=0 2x = 5 x = 5/2 which means that x = 5/2 makes the denominator zero. So to avoid dividing by zero, we just add the restriction that x cannot be 5/2 So the answer is choice B

OpenStudy (anonymous):

thank you

jimthompson5910 (jim_thompson5910):

you're welcome